Circular No. 7651

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cx IN UGC 12266 C. Yu, M. Modjaz, and W. D. Li, University of California at Berkeley,

report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.1) in an unfiltered image taken with the Katzman Automatic Imaging Telescope (KAIT) on June 25.5 UT (confirmed on an earlier image taken on June 9.5 at mag ∼ 17.9 under partly cloudy conditions). The new object is located at α = 22^{h}58^{m}00^{s}.28, δ = +6^{o}03′48′′.4 (equinox 2000.0), which is 2′′.9 west and 23′′.3 south of the nucleus of UGC 12266. A KAIT image taken on 2000 Dec. 12.3 showed nothing at this position (limiting mag ∼ 19.0).

V1178 SCORPII M. Fujii, Okayama, Japan, reports that a spectrum (resolution ∼ 1 nm;

instrumental resolution ∼ 450 km/s) of V1178 Sco (cf. IAUC 7647, 7650), obtained on June 24.57 UT with a 0.28-m reflector, shows a very strong Hα emission line (FWHM ∼ 1300 km/s) and weaker Hβ. The signal-to-noise ratio was rather low due to clouds.

β CEPHEI

C. Neiner, Observatoire de Meudon and University of Amsterdam (UA); H. Henrichs and V. Geers, UA; and J.-F. Donati, Observatoire de Midi-Pyrénées, Toulouse, write: “The slowly rotating, pulsating, magnetic star β Cep (spectral type B1 IVe) has again gone into an Hα emission phase. In June 1999, Hα was still completely in absorption, with emission nearly absent. On 2001 June 19, observations made with the Pic du Midi 2-m telescope (+ MuSiCoS spectrograph; resolution 38 000) show the doublepeaked emission (with a stronger red peak) extended above the continuum from −400 to +400 km/s. The previous Hα outburst occurred in 1990.”

SUPERNOVA 2001cj IN UGC 8399 CCD photometry, unfiltered unless otherwise noted: June 7.129 UT,

15.77 (J. D. West, Mulvane, KS); 9.138, 15.78 (West); 10.143, 15.85 (West); 12.138, 15.83 (West); 12.979, R = 16.08 (K. Hornoch, Lelekovice, Czech Republic); 13.887, R = 16.13 (Hornoch); 14.154, 16.4 (West); 15.800, R = 16.17 (Hornoch); 20.925, R = 16.37 (Hornoch); 21.869, 16.2 (G. Masi, Ceccano, Italy); 24.899, R = 16.56 (Hornoch); 25.977, R = 16.67 (Hornoch).

2001 June 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green